imloyalipromise:

Right triangle ABC is located at A (−1, −2), B (−1, 1), and C (−5, 1) on a coordinate plane. What is the equation of a circle A with radius segment AC? (x + 1)2 + (y + 2)2 = 9 (x + 5)2 + (y − 1)2 = 16 (x + 1)2 + (y + 2)2 = 25 (x + 5)2 + (y − 1)2 = 25

Aratox:

Do you have an online/paper representation of the graph? If you do, please post it

Darkest:

Uum I ain't done this in a little minute lemme think

imloyalipromise:

@aratox wrote:
Do you have an online/paper representation of the graph? If you do, please post it
there is no graph its just that

Darkest:

C

hannahj:

(x +1)^2 +(y +2)^2 = 25

imloyalipromise:

@hannahj wrote:
(x +1)^2 +(y +2)^2 = 25
thanks

HotPockets14:

@hannahj wrote:
(x +1)^2 +(y +2)^2 = 25
this is the correct answer

Luigi0210:

@hannahj wrote:
(x +1)^2 +(y +2)^2 = 25
And where's the explanation..?

hannahj:

The center of the circle is located at point A, which is (-1,-2). The radius of the circle is the length of segment AC. First, find the length of the radius by calculating the distance between A and C using the distance formula: √[(x2-x1)² + (y2-y1)²] = √[(-5 - (-1))² + (1 - (-2))²] = √[(-4)² + 3²] = √[16 + 9] = √25 = 5. Then, the equation of the circle is (x-h)² + (y-k)² = r², where (h,k) is the center and r is the radius. Substituting (-1,-2) for (h,k) and 5 for r, the equation becomes (x+1)² + (y+2)² = 5², or (x + 1)² + (y + 2)² = 25. So, the answer is option C.

L8RAaronOnTop:

W